kamachavis:

h(x)=x^2-9/x^2+4 Identify the function, can fall into several categories.

jhonyy9:

\[h(x) = \frac{ x^2 -9 }{ x^2 +4 }\] in this way wrote ? @supie

kamachavis:

Yes.

jhonyy9:

so how you think when there is a fractional equation with variable on denominator what is the first step what you need to do ? by what not can be equal the denominator never ?

kamachavis:

0

jhonyy9:

so remember a fraction to not be yes - good

jhonyy9:

so for what value of x you get the denominator zero ?

jhonyy9:

bc. the denominator has a x^2 term so this not will result a negativ number never yes ?

kamachavis:

yes

jhonyy9:

ok and now what you see what is in the numerator of this fraction ?

jhonyy9:

x^2 -9

kamachavis:

yes

jhonyy9:

x^2 -3^2 what is this ?

jhonyy9:

do you know formula of ?

kamachavis:

no

Extrinix:

its been so long since ive done this-

jhonyy9:

difference of two squares

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Extrinix its been so long since ive done this- \(\color{#0cbb34}{\text{End of Quote}}\) please dont interfere never - the rule of QC

jhonyy9:

@Laylalyssa

jhonyy9:

please explain

kamachavis:

answer choices vary from rational, algebraic, polynomial degree of 2 and some answers consist of multiple of those choices

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kamachavis answer choices vary from rational, algebraic, polynomial degree of 2 and some answers consist of multiple of those choices \(\color{#0cbb34}{\text{End of Quote}}\) ok and this is true that was solved by @Extrinix ?

kamachavis:

no

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Extrinix its been so long since ive done this- \(\color{#0cbb34}{\text{End of Quote}}\) why he said this ?

jhonyy9:

@Extrinix ???

Extrinix:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @Extrinix ??? \(\color{#0cbb34}{\text{End of Quote}}\) hmm?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Extrinix its been so long since ive done this- \(\color{#0cbb34}{\text{End of Quote}}\) why he said this ? \(\color{#0cbb34}{\text{End of Quote}}\) you wrote this above - why ?

jhonyy9:

you ve done this - when ?

Extrinix:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 you ve done this - when ? \(\color{#0cbb34}{\text{End of Quote}}\) ive done similar equations like this a while back... its just been so long its hard to remember exactly how i did it yk

jhonyy9:

@supie any idea ?

jhonyy9:

x^2 -9 = (x-3)(x+3)

jhonyy9:

who asked go away - is offline